OpenStudy (anonymous):

Factor completely x2 – 12x + 35

OpenStudy (unklerhaukus):

\[x^2-12x+35\]\[=x^2-ax-bx+35\]\[=(x-a)(x-b)\] \[a\times b=35\]

OpenStudy (unklerhaukus):

\[a+b=12\]

OpenStudy (anonymous):

thnx

OpenStudy (unklerhaukus):

have you found a,b?